This Question: 4 pts Evaluate the integral.
Answer #1

It brettas - ( 165-ci se I - (55-65lads I = (55-65+ (as+byh lastb) ntl ( 31+) •f. (55-6+C 1 = ds 1 (n+1)•7 (astb] as (55-63 1